# Statistical reduced order modelling for the parametric Helmholtz equation

### v1

@@ 4.1 Forward Problem
The statFEM approach [4] is in its core based on Gaussian Processes (GPs), which form a class of Bayesian non-parametric models that are well established in statistics and machine learning [23]. In a first step, a GP model over the FE solution is formed, which is obtained by propagating uncertainty in material parameters, the geometry or excitations through the FE model. Here, to account for the uncertainty both in the material parameter and the right-hand-side, a ROM is used. Then, in a second step, data can be used for updating the state and for reducing the prediction uncertainty. In this way, statFEM makes use of GPs to fuse a solution from a FE solver with data. When modeling the state as a GP, normally distributed model and data errors give rise to a normally distributed posterior, hence Bayesian updating can be carried out in closed form. This in turn is convenient for a monitoring perspective, where fast sampling free updating procedures are desired.
@@
$c_{g}(\bm{x},\bm{x}^{\prime})=\sigma^{2}\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\|\bm{x}-\bm{x}^{\prime}\|}{l}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\|\bm{x}-\bm{x}^{\prime}\|}{l}\right),$ (23)
with the scaling parameter $\sigma\in\mathbb{R}^{+}$, $\Gamma(\nu)$ the Gamma function and $K_{\nu}$ a modified Bessel function [24, p.84 ff.], [5]. Moreover, $l\in\mathbb{R}^{+}$ is the lengthscale parameter and $\nu\in\mathbb{R}^{+}$ controls the smoothness of the kernel. For $\nu\to\infty$ (23) becomes the squared exponential kernel. The Matérn kernel is used here because of its flexibility to represent variable smoothness.
@@
$(\bm{C}_{g})_{i,j}=\left(\int_{\Gamma_{\mathrm{N},g}}\phi_{i}\,\mathrm{d}x\right)c_{g}(\bm{x}_{i},\bm{x}_{j})\left(\int_{\Gamma_{\mathrm{N},g}}\phi_{j}^{*}\,\mathrm{d}x\right).$ (25)
In view of the linear dependence of $\bm{f}_{g}$ on $g$, the source term can be represented by a multivariate Gaussian
@@
We also assume that the remaining stochastic forcing over the domain, which by a slight abuse of notation is again denoted with $f$, is a Gaussian random field
@@
with $c_{f}(\bm{x},\bm{x}^{\prime})$ defined again via the Matérn kernel. Applying the same discretization principle as before, we obtain
@@
$\bm{A}(\omega,\bm{\xi}(\theta))\bm{u}(\theta)=\bm{f}(\theta)+\bm{f}_{g}(\theta),$ (30)
where $\bm{\xi},\bm{f},\bm{f}_{g}$ are independent random vectors on the same probability space.
In the original statFEM paper [4], the authors employ a perturbation approach to propagate forward probabilistic material coefficients that appear in the FEM system matrix. We, instead, apply a Quasi Monte Carlo (QMC) method [26] with the joint input $\bm{\eta}=(\bm{\xi}^{\top},\bm{f}^{\top},\bm{f}_{g}^{\top})^{\top}$ to approximate the prior mean and covariance of $\bm{u}$. To this end, let $\bm{\eta}^{(i)}$, denote a sequence based on digital nets with random shifting [27], where the inverse CDF is used to rescale the QMC points initially defined on $[0,1]$. Note that, because of the linearity of $\bm{A}$, the randomness in the right-hand-side could be propagated analytically to the solution. However, we employ the QMC sampling approach in the same way for all uncertain inputs, for simplicity. We construct for each sample $\bm{\xi}^{(i)}$ the ROM projection matrix $\bm{V}^{(i)}$ as defined in Section 3. Then, we can efficiently solve the system
$(\bm{V}^{(i)})^{*}\bm{A}(\omega,\bm{\xi}^{(i)})\bm{V}^{(i)}\bm{u}^{(i)}(\omega)=(\bm{V}^{(i)})^{*}(\bm{f}^{(i)}+\bm{f}_{g}^{(i)}),$ (31)
@@
After applying the standard QMC-estimates of the mean and covariance function based on the sample points $\bm{u}^{(i)}$, the final ROM-QMC approximated, parameter-dependent, GP prior reads
$\bm{u}(\omega)\approx\tilde{\bm{u}}(\omega)\sim\mathcal{N}(\bm{\mu}_{\tilde{\bm{u}}}(\omega),\bm{C}_{\tilde{\bm{u}}}(\omega)).$ (32)
@@
In (33), the data $\bm{y}\in\mathbb{R}^{n_{y}}$ are decomposed into the true but unknown states at the $n_{y}$ sensor locations $\bm{z}\in\mathbb{R}^{n_{y}}$ and the data error $\bm{e}\in\mathbb{R}^{n_{y}}$. The true state $\bm{z}$ can further be expressed as a combination of the FE prediction $\bm{u}\in\mathbb{R}^{N}$, scaled by an unknown factor $\rho$ and projected onto the sensor locations with the projection matrix $\bm{P}\in\mathbb{R}^{n_{y}\times N}$, and a statistical representation of the model error $\bm{d}\in\mathbb{R}^{n_{y}}$.
For the time being, we assume that the prior $p(\bm{u})\sim\mathcal{N}(\bm{\mu}_{\bm{u}},\bm{C}_{\bm{u}})$ is fully available and postpone the discussion on numerical approximations errors. Then, the model error is given the difference between the exact model response and the true state at the sensor locations and essentially a deterministic quantity, that can be reduced by improving the model. The choice of modeling $\bm{d}$ in a statistical way is made out of convenience, but also to improve the data assimilation accuracy, see [29] for example. More precisely, following [4] we set
@@
with $\bm{\alpha}$ the hyperparameters of the chosen kernel function. Here, the Matérn kernel (23) is also chosen for the model error. The hyperparameters are estimated by maximising the marginal likelihood, i.e. the likelihood of the data given the hyperparameters $p(\bm{\bm{y}|\bm{\alpha}})$. This is detailed in Section 4.4. Please note, that the scale of the model error will be reflected in the parameter $\sigma$ of (23). Hence, an improvement of the model will result in a smaller estimated value of $\sigma$ and hence a smaller uncertainty band for the model error, see also [30]. Evaluating the kernel function at the measurement positions, the GP is represented by a multivariate Gaussian with a covariance matrix $\bm{C_{d}}\in\mathbb{R}^{n_{y}\times n_{y}}$.
@@
Finally, the predicted data distribution at new sensor locations, with $\hat{\bm{P}}\in\mathbb{R}^{n_{\hat{y}}\times n_{u}}$, reads
$p(\hat{\bm{y}}|\bm{Y})=\mathcal{N}(\rho\hat{\bm{P}}\bm{\mu}_{\bm{u}|\bm{y}},\rho^{2}\bm{P}\bm{C}_{\bm{u}|\bm{y}}\hat{\bm{P}}^{T}+\hat{\bm{C}}_{d}+\hat{\bm{C}}_{e})$ (39)
@@ 4.3 ROM error estimation
We assume that the error is smooth enough over the computational domain so that again a GP model can be used to represent it. In particular, we estimate the ROM error at a few randomly chosen points and condition a GP to obtain an approximation over the entire domain. It is important to quantify the ROM error over the entire domain, not only at the sensor locations, to improve also the prediction of the true response $\bm{z}|\bm{Y}$ and $\hat{\bm{y}}|\bm{Y}$. Since, in this approach, the error is only needed at a few points, we can employ adjoint methods. We follow an approach similar to the one presented in [15], where the authors estimated the full FEM error field, contrary to the full ROM error field as desired here. Consider the solution $\bm{u}^{(i)}$ and the adjoint equation
@@
where $\bm{p}_{\ell}^{*}\bm{u}(\omega)\approx u(\omega,\bm{x}_{\ell})$. Then, the error related to $\bm{x}_{\ell}$ can be rewritten as
@@
$=((\bm{A}(\omega,\bm{\xi}^{(i)}))^{*}\bm{q}_{\ell}^{(i)}(\omega))^{*}\left(\bm{u}^{(i)}(\omega)-\bm{V}^{(i)}\bm{u}_{r}^{(i)}(\omega)\right)$ (42)
@@
$=(\bm{q}_{\ell}^{(i)}(\omega))^{*}\left(\bm{f}^{(i)}+\bm{f}_{g}^{(i)}-\bm{A}(\omega,\bm{\xi}^{(i)})\bm{V}^{(i)}\bm{u}_{r}^{(i)}(\omega)\right).$ (44)
@@
where $(\bm{A}_{r}(\omega,\bm{\xi}^{(i)}))^{*}=(\bm{V}_{\mathrm{adj},\ell}^{(i)})^{*}(\bm{A}(\omega,\bm{\xi}^{(i)}))^{*}\bm{V}_{\mathrm{adj},\ell}^{(i)}$, $\bm{q}_{r,\ell}^{(i)}(\omega)=\bm{V}_{\mathrm{adj},\ell}^{(i)}\bm{q}_{\ell}^{(i)}(\omega)$, $\bm{p}_{r,\ell}^{(i)}=(\bm{V}_{\mathrm{adj},\ell}^{(i)})^{*}\bm{p}_{\ell}$ and $\bm{q}_{\ell}^{(i)}(\omega)=\bm{V}_{\mathrm{adj},\ell}^{(i)}\bm{q}_{r,\ell}^{(i)}(\omega)$. Please note that the construction of $\bm{V}_{\mathrm{adj},\ell}^{(i)}\neq\bm{V}^{(i)}$ has to be carried out additionally in the offline phase. However, running the moment matching algorithm again for $i=1,\ldots,I$ is very cheap as we can reuse the LU decomposition of $\bm{A}(\omega,\bm{\xi}^{(i)})$ which had to be computed already to find the ROM prior.
Remark.
@@
In the next sections, this ROM approximation error indicator will be included into the statFEM posterior update and it will be shown that this can help improving the posterior mean estimate.
Remark.
If the problem is subject only to a stochastic right hand side - called $\bm{f}$ for simplicity - it is possible to compute a prior for $\bm{d}_{r}$ in closed form. In turn, the number of required projection matrices can be reduced. Let $\bm{f}_{r}\sim\mathcal{N}(\bm{\mu}_{\bm{f}_{r}},\bm{C}_{\bm{f}_{r}})$ be the reduced stochastic right-hand-side. We propagate its distribution to $\bm{u}_{r}$ via
@@
$\bm{q}_{\ell}^{*}\bm{A}(\omega)\bm{V}\bm{u}_{r}(\omega)\sim\mathcal{N}(\bm{q}_{\ell}^{*}\bm{B}(\omega)\bm{\mu}_{\bm{f}_{r}},(\bm{q}_{\ell}^{*}\bm{B}(\omega))^{*}\bm{V}^{*}\bm{C}_{f}\bm{V}(\bm{q}_{\ell}^{*}\bm{B}(\omega)))$ (52)
with $\bm{B}=\bm{A}\bm{V}\bm{A}_{r}^{-1}$. Please note that $\bm{V}$ is constructed with a deterministic $\bm{A}$ and a specific realization of $\bm{f}$ as before - and hence, is deterministic in this case. Therefore, we obtain the following prior for the ROM-error
@@
A standard statFEM posterior update can of course be conducted, with the prior FE solution replaced with a ROM. However, the error model employed in (55) is rather crude and the additional ROM approximation error would simply inflate $\bm{C}_{\bm{d}}$. In the last section, we have formulated a numerically inexpensive way to approximate the ROM error for every point in the domain. Knowing more about the numerical error should, intuitively, lead to a better solution, i.e., a sharper estimated model error and, for the predictive observation density, a posterior mean closer to the observed data. Therefore, in this section, the statistical generating model and the posterior update equations are modified to account for the ROM approximation error. We label the resulting updating procedure as statistical reduced order modeling (statROM).
@@
with $\bm{d}_{r}\in\mathbb{R}^{N}$ representing the ROM approximation error, which is obtained by interpolating the ROM error GP at the FE mesh nodes. Unlike the model error $\bm{d}$, the ROM error indicator involves a non-zero mean. Hence, a bias induced by the ROM can directly be corrected and the ROM approximation error does not only inflate the variance. The statistical generating model is now used to derive the posterior update equations.
@@
Conditioning $\tilde{\bm{u}}$ on $\bm{y}$ now yields for the posterior mean
$\begin{split}\bm{\mu}_{\tilde{\bm{u}}|\bm{y}}=\bm{\mu}_{\tilde{\bm{u}}}+\rho\bm{C}_{\tilde{u}}\bm{P}^{\top}\left[\rho^{2}\bm{P}\bm{C}_{\tilde{\bm{u}}}\bm{P}^{\top}+\bm{K}_{\bm{y},r}\right]^{-1}\left(\bm{y}-\left(\rho\bm{P}\bm{\mu}_{\tilde{\bm{u}}}+\bm{P}\bm{\mu}_{\bm{d}_{r}}\right)\right)\end{split}$ (57)
with with $\bm{K}_{\bm{y}_{r}}=\rho^{2}\bm{P}\bm{C}_{\bm{d}_{r}}\bm{P}^{\top}+\bm{C}_{\bm{d}}+\bm{C}_{\bm{e}}$ and for the covariance
@@
where we recall that $\tilde{\bm{u}}\sim\mathcal{N}(\bm{\mu}_{\tilde{\bm{u}}},\bm{C}_{\tilde{\bm{u}}})$ is the ROM-QMC prior solution. For multiple readings, i.e. $n_{o}>1$, we obtain
@@
Hence, for the statROM posterior there holds
@@
The equations can also be derived by directly using Bayes’ law as shown in [4]. Adapting the argumentation of [4] to the present setting, we can deduce that for a covariance term $\bm{K}_{\bm{y},r}$ which is small in comparison to $\bm{C}_{\tilde{\bm{u}}}$, the posterior mean tends to $(\sum_{i=1}^{n_{o}}\bm{y}_{i}-n_{o}\bm{\mu}_{\bm{d}_{r}})/(\rho n_{o})$ and the posterior covariance tends to $\bm{K}_{\bm{y}_{r}}/\rho^{2}$. For a large covariance term, the posterior is close to the prior. The higher the number of observations $n_{o}$, the more the posterior covariance tends to zero and the mean to the empirical mean of the ROM-error corrected observations $(\sum_{i=1}^{n_{o}}\bm{y}_{i}-n_{o}\bm{\mu}_{\bm{d}_{r}})/n_{o}$.

### v2

@@ 4.1 Forward Problem
The statFEM approach [4] is in its core based on GPs, which form a class of Bayesian non-parametric models that are well established in statistics and machine learning [34]. In a first step, a GP model over the FE solution is formed, which is obtained by propagating uncertainty in material parameters, the geometry or excitations through the FE model. Here, to account for the uncertainty both in the material parameter and the right-hand-side, a ROM is used. Then, in a second step, data can be used for updating the state and for reducing the prediction uncertainty. In this way, statFEM makes use of GPs to fuse a solution from a FE solver with data. When modelling the state as a GP, normally distributed model and data errors give rise to a normally distributed posterior, hence Bayesian updating can be carried out in closed form. This in turn is convenient for a monitoring perspective, where fast sampling free updating procedures are desired.
@@
[…]
with the scaling parameter $\sigma\in\mathbb{R}^{+}$, $\Gamma(\nu)$ the Gamma function and $K_{\nu}$ a modified Bessel function [35, p.84 ff.], [5]. Moreover, $l\in\mathbb{R}^{+}$ is the lengthscale parameter and $\nu\in\mathbb{R}^{+}$ controls the smoothness of the kernel. For $\nu\to\infty$ (22) becomes the squared exponential kernel. The Matérn kernel is used here because of its flexibility to represent variable smoothness.
@@
$(\bm{C}_{g})_{i,j}=\int_{\Gamma_{\mathrm{N},g}}\int_{\Gamma_{\mathrm{N},g}}\phi_{i}(\bm{x})c_{g}(\bm{x},\bm{x}^{\prime})\phi_{j}^{*}(\bm{x}^{\prime})\,\mathrm{d}x\,\mathrm{d}x^{\prime}\\ \approx\left(\int_{\Gamma_{\mathrm{N},g}}\phi_{i}\,\mathrm{d}x\right)c_{g}(\bm{x}_{i},\bm{x}_{j})\left(\int_{\Gamma_{\mathrm{N},g}}\phi_{j}^{*}\,\mathrm{d}x\right).$ (24)
In (24), we follow [4] and interpolate the covariance function with the FE basis functions. Mass lumping then leads to the approximation in (24) that still will be denoted as $(\bm{C}_{g})_{i,j}$ in the following. The associated approximation error is expected to be small, at the same time, the implementation largely simplifies. We refer to [4] and the references therein for a discussion. In view of the linear dependence of $\bm{f}_{g}$ on $g$, the source term can be represented by a multivariate Gaussian
@@
In addition to $\bm{f}_{g}$, we consider a volume source term on the right-hand-side, corresponding to $f$ in (1). This term is model as a Gaussian random field
@@
as well, with $c_{f}(\bm{x},\bm{x}^{\prime})$ defined via the Matérn kernel. Applying the same discretization principle as before, cf. (23) - (25), we obtain
@@
$\bm{A}(\omega,\bm{\xi}(\theta))\bm{u}(\omega,\theta)=\bm{F}(\omega,\theta),$ (29)
where $\bm{F}(\omega,\theta)=\bm{f}(\omega,\theta)+\bm{f}_{g}(\omega,\theta)$ and $\bm{\xi},\bm{f}(\omega,\cdot),\bm{f}_{g}(\omega,\cdot)$ are independent random vectors on the same probability space, collected as $\bm{\eta}=(\bm{\xi}^{\top},\bm{f}(\omega,\cdot)^{\top},\bm{f}_{g}(\omega,\cdot)^{\top})^{\top}\in\mathbb{R}^{M}\times\mathbb{C}^{2N}$.
In the original statFEM paper [4], the authors employ a perturbation approach to propagate forward probabilistic material coefficients that appear in the FEM system matrix. We, instead, apply a quasi Monte Carlo (QMC) method [15] to approximate the prior mean and covariance of $\bm{u}$. To this end, let $\tilde{\bm{\eta}}^{(i)}\in[0,1]^{M+2N}\ i=1,\ldots,Q$, denote a sequence based on digital nets with random shifting [37]. Since all elements of $\bm{\eta}$ are normally distributed, we use the transformation
$\bm{\eta}^{(i)}=\left(\begin{array}[]{c}\bm{\xi}^{(i)}\\ \bm{f}^{(i)}(\omega)\\ \bm{f}_{g}^{(i)}(\omega)\end{array}\right)=\left(\begin{array}[]{c}\bm{0}\\ \bm{\mu}_{f}(\omega)\\ \bm{\mu}_{g}(\omega)\end{array}\right)+\left(\begin{array}[]{ccc}\bm{I}&\bm{0}&\bm{0}\\ \bm{0}&\bm{C}_{f}^{1/2}&\bm{0}\\ \bm{0}&\bm{0}&\bm{C}_{g}^{1/2}\end{array}\right)\Phi^{-1}(\tilde{\bm{\eta}}^{(i)})$ (30)
where $\Phi$ denotes the CDF of a standard Gaussian random variable, and the inverse is applied component-wise in (30). Note that, because of the linearity of $\bm{A}$, the randomness in the right-hand-side could be propagated analytically to the solution. However, we employ the QMC sampling approach in the same way for all uncertain inputs, for simplicity. We construct for each sample $\bm{\xi}^{(i)}$ the ROM projection matrix $\bm{V}^{(i)}$ as defined in Section 3. Then, we can efficiently solve the system
$(\bm{V}^{(i)})^{*}\bm{A}(\omega,\bm{\xi}^{(i)})\bm{V}^{(i)}\bm{u}_{r}^{(i)}(\omega)=(\bm{V}^{(i)})^{*}\bm{F}^{(i)}(\omega),$ (31)
@@
Given the the ROM projection matrices $\bm{V}^{(i)},i=1,\ldots,Q$, for every query point $\omega$, we compute a Gaussian approximation as
$\bm{u}(\omega)\approx\bm{u}_{Q,r}(\omega)\sim\mathcal{N}(\bm{\mu}_{\bm{u};Q,r}(\omega),\bm{C}_{\bm{u};Q,r}(\omega)),$ (32)
where we first solve for $\bm{u}_{r}^{(i)}(\omega),i=1,\ldots,Q$ using (31) and then build the standard Monte Carlo mean and covariance approximation
$\bm{\mu}_{\bm{u}}(\omega)$ $\approx\bm{\mu}_{\bm{u};Q,r}(\omega)=\frac{1}{Q}\sum_{i=1}^{Q}\bm{V}^{(i)}\bm{u}^{(i)}(\omega)$ (33)
$\bm{C}_{\bm{u}}(\omega)$ $\approx\bm{C}_{\bm{u};Q,r}(\omega)$
$=\frac{1}{Q-1}\sum_{i=1}^{Q}\left(\bm{V}^{(i)}\bm{u}^{(i)}(\omega)-\bm{\mu}_{\bm{u};Q,r}(\omega)\right)\left(\bm{V}^{(i)}\bm{u}^{(i)}(\omega)-\bm{\mu}_{\bm{u};Q,r}(\omega)\right)^{*}.$ (34)
A sample for $\bm{u}_{Q,r}(\omega)$ is then generated in the standard way as $\bm{\mu}_{\bm{u};Q,r}(\omega)+\bm{C}_{\bm{u};Q,r}^{1/2}(\omega)\bm{\chi}$, with $\bm{\chi}\sim\mathcal{N}(\bm{0},\bm{I})$ of dimension $N$. In the following, to simplify the notation, we will denote the QMC-ROM Gaussian approximation as $\tilde{\bm{u}}=\bm{u}_{Q,r}$ and drop the dependence on $\omega$ in the next subsection.
@@
In (35), the data $\bm{y}\in\mathbb{C}^{n_{y}}$ are decomposed into the true but unknown states $\bm{z}\in\mathbb{C}^{n_{y}}$, at the sensor locations $\check{\bm{x}}_{i},i=1,\ldots,n_{y}$ and the data error $\bm{e}\in\mathbb{C}^{n_{y}}$. The true state $\bm{z}$ can further be expressed as a combination of the FE prediction $\bm{u}\in\mathbb{C}^{N}$, scaled by an unknown factor $\rho$ and projected onto the sensor locations with the projection matrix $\bm{P}\in\mathbb{R}^{n_{y}\times N}$, and a statistical representation of the model error $\bm{d}\in\mathbb{C}^{n_{y}}$. According to [4], the projection matrix $\bm{P}$ is computed by evaluating the FE basis functions at the sensor locations. Hence, to use statFEM with commercial FE codes, these at least need to provide information about the used basis functions or provide them explicitly.
For the time being, we assume that the prior $p(\bm{u})\sim\mathcal{N}(\bm{\mu}_{\bm{u}},\bm{C}_{\bm{u}})$ is fully available and postpone the discussion about numerical approximation errors. Then, the model discrepancy is given by the difference between the exact model response and the true state at the sensor locations. This is essentially a deterministic quantity, that can be reduced by improving the model. The choice of modelling $\bm{d}$ in a statistical way is made out of convenience, but also to improve the data assimilation accuracy, see [39] for example. More precisely, following [4] we set
@@
[…]
$(\bm{C_{\bm{d}}}(\bm{\alpha}))_{ij}=c_{d}(\check{\bm{x}}_{i},\check{\bm{x}}_{j};\bm{\alpha}).$ (37)
The hyperparameters are estimated by maximising the marginal likelihood, i.e. the likelihood of the data given the hyperparameters $p(\bm{\bm{y}|\bm{\alpha}})$. This is detailed in Section 4.4. Please note, that the scale of the model error will be reflected in the parameter $\sigma$ of (22). Hence, an improvement of the model will result in a smaller estimated value of $\sigma$ and hence a smaller uncertainty band for the model error, see also [40]. Evaluating the kernel function at the measurement positions, the GP is represented by a multivariate Gaussian with a covariance matrix $\bm{C_{d}}\in\mathbb{R}^{n_{y}\times n_{y}}$.
@@
Finally, the predicted data distribution at new sensor locations, with $\hat{\bm{P}}\in\mathbb{R}^{n_{\hat{y}}\times N}$, reads
$p(\hat{\bm{y}}|\bm{Y})=\mathcal{N}(\rho\hat{\bm{P}}\bm{\mu}_{\bm{u}|\bm{Y}},\rho^{2}\hat{\bm{P}}\bm{C}_{\bm{u}|\bm{Y}}\hat{\bm{P}}^{T}+\hat{\bm{C}}_{d}+\hat{\bm{C}}_{e})$ (42)
@@ 4.3 ROM error estimation
We assume that the error is smooth enough over the computational domain so that again a GP model can be used to represent it. In particular, we estimate the ROM error at a few randomly chosen points and condition a GP to obtain an approximation over the entire domain. It is important to quantify the ROM error over the entire domain, not only at the sensor locations, to improve also the prediction of the true response $\bm{z}|\bm{Y}$ and $\hat{\bm{y}}|\bm{Y}$. Since, in this approach, the error is only needed at a few ($n_{d}$) points, we can employ adjoint methods. We follow an approach similar to the one presented in [17], where the authors estimated the full FEM error field, contrary to the full ROM error field as desired here. Consider the solution $\bm{u}^{(i)}$ and the adjoint equation
@@
where $l=1,\ldots,n_{d}$ and $\bm{p}_{\ell}^{*}\bm{u}(\omega)\approx u(\omega,\bm{x}_{\ell})$. Then, the error related to $\bm{x}_{\ell}$ can be rewritten as
@@
$=\left((\bm{A}(\omega,\bm{\xi}^{(i)}))^{*}\bm{q}_{\ell}^{(i)}(\omega)\right)^{*}\left(\bm{u}^{(i)}(\omega)-\bm{V}^{(i)}\bm{u}_{r}^{(i)}(\omega)\right)$ (45)
@@
$=(\bm{q}_{\ell}^{(i)}(\omega))^{*}\left(\bm{F}^{(i)}(\omega)-\bm{A}(\omega,\bm{\xi}^{(i)})\bm{V}^{(i)}\bm{u}_{r}^{(i)}(\omega)\right).$ (47)
@@
where $(\bm{A}_{r}(\omega,\bm{\xi}^{(i)}))^{*}=(\bm{V}_{\mathrm{adj},\ell}^{(i)})^{*}(\bm{A}(\omega,\bm{\xi}^{(i)}))^{*}\bm{V}_{\mathrm{adj},\ell}^{(i)}$, $\bm{p}_{r,\ell}^{(i)}=(\bm{V}_{\mathrm{adj},\ell}^{(i)})^{*}\bm{p}_{\ell}$ and $\bm{q}_{\ell}^{(i)}(\omega)=\bm{V}_{\mathrm{adj},\ell}^{(i)}\bm{q}_{r,\ell}^{(i)}(\omega)$. Please note that the construction of $\bm{V}_{\mathrm{adj},\ell}^{(i)}\neq\bm{V}^{(i)}$ has to be carried out additionally in the offline phase. However, running the moment matching algorithm again for $i=1,\ldots,I$ is very cheap as we can reuse the LU decomposition of $\bm{A}(\omega,\bm{\xi}^{(i)})$ which had to be computed already to find the ROM prior.
Remark 2.
@@
In the next sections, this ROM approximation error indicator will be included into the statFEM posterior update and it will be shown that this can help improve the posterior mean estimate.
Remark 3.
If the problem is subject only to a stochastic right hand side $\bm{F}$ it is possible to compute a prior for $\bm{d}_{r}$ in closed form. In turn, the number of required projection matrices can be reduced. Let $\bm{F}_{r}\sim\mathcal{N}(\bm{\mu}_{\bm{F}_{r}},\bm{C}_{{\bm{F}}_{r}})$ be the reduced stochastic right-hand-side. We propagate its distribution to $\bm{u}_{r}$ via
@@
$\bm{q}_{\ell}^{*}\bm{A}(\omega)\bm{V}\bm{u}_{r}(\omega)\sim\mathcal{N}(\bm{q}_{\ell}^{*}\bm{B}(\omega)\bm{\mu}_{{\bm{F}}_{r}},(\bm{q}_{\ell}^{*}\bm{B}(\omega))^{*}\bm{V}^{*}\bm{C}_{\bm{F}}\bm{V}(\bm{q}_{\ell}^{*}\bm{B}(\omega)))$ (55)
with $\bm{B}=\bm{A}\bm{V}\bm{A}_{r}^{-1}$. Please note that $\bm{V}$ is constructed with a deterministic $\bm{A}$ and a specific realization of $\bm{F}$ - as before - and hence, is deterministic in this case. Therefore, we obtain the following prior for the ROM-error
@@
A standard statFEM posterior update can of course be conducted, with the prior FE solution replaced with a ROM. However, the error model employed in (35) is rather crude and the additional ROM approximation error would simply inflate $\bm{C}_{\bm{d}}$. In the last section, we have formulated a way to approximate the ROM error for every point in the domain. Knowing more about the numerical error should, intuitively, lead to a better solution, i.e., a sharper estimated model error and, for the predictive observation density, a posterior mean closer to the observed data. Therefore, in this section, the statistical generating model and the posterior update equations are modified to account for the ROM approximation error. We label the resulting updating procedure as statistical reduced order modelling (statROM).
@@
with $\bm{d}_{r}\in\mathbb{C}^{N}$ representing the ROM approximation error, which is obtained by interpolating the ROM error GP at the FE mesh nodes. Unlike the model error $\bm{d}$, the ROM error indicator involves a non-zero mean. Hence, a bias induced by the ROM can directly be corrected and the ROM approximation error does not only inflate the variance. The statistical generating model is now used to derive the posterior update equations.
@@
Conditioning $\tilde{\bm{u}}$ on $\bm{y}$ now yields, using the Sherman-Morris-Woodbury identity, for the posterior mean
$\begin{split}\bm{\mu}_{\tilde{\bm{u}}|\bm{y}}=\bm{\mu}_{\tilde{\bm{u}}}+\rho\bm{C}_{\tilde{u}}\bm{P}^{\top}\left[\rho^{2}\bm{P}\bm{C}_{\tilde{\bm{u}}}\bm{P}^{\top}+\bm{K}_{\bm{y}_{r}}\right]^{-1}\left(\bm{y}-\left(\rho\bm{P}\bm{\mu}_{\tilde{\bm{u}}}+\rho\bm{P}\bm{\mu}_{\bm{d}_{r}}\right)\right)\end{split}$ (60)
with $\bm{K}_{\bm{y}_{r}}=\rho^{2}\bm{P}\bm{C}_{\bm{d}_{r}}\bm{P}^{\top}+\bm{C}_{\bm{d}}+\bm{C}_{\bm{e}}$ and for the covariance
@@
[…]
@@
Hence, for the statROM posterior we have
@@
The equations can also be derived by directly using Bayes’ law as shown in [4]. Adapting the argumentation of [4] to the present setting, we can deduce that for a covariance term $\bm{K}_{\bm{y}_{r}}$ which is small in comparison to $\bm{C}_{\tilde{\bm{u}}}$, the posterior mean tends to $(\sum_{i=1}^{n_{o}}\bm{y}_{i}-n_{o}\bm{\mu}_{\bm{d}_{r}})/(\rho n_{o})$ and the posterior covariance tends to $\bm{K}_{\bm{y}_{r}}/\rho^{2}$. For a large covariance term, the posterior is close to the prior. The higher the number of observations $n_{o}$, the more the posterior covariance tends to zero and the mean to the empirical mean of the ROM-error corrected observations $(\sum_{i=1}^{n_{o}}\bm{y}_{i}-n_{o}\bm{\mu}_{\bm{d}_{r}})/n_{o}$.